A force F=100N points 300 with respects to the x-axis. The force acts on an object of mass 5kg.

a. what is the magnitude and direction of the acceleration of that object ? [3+2 marks]

b. what are the magnitudes of the x-components of the force and the acceleration (Fx and ax)?


1
Expert's answer
2022-03-15T10:35:39-0400

(a)

"a=\frac{F}{m}=\frac{100\:\rm N}{5\: \rm kg}=20\:\rm m/s^2"

Acceleration directed at angle "30^\circ" with respects to the x-axis.

(b)

"F_x=F\cos\theta=\rm 100\: N\cos 30^\circ=86.6\: N"

"a_x=\frac{F_x}{m}=\frac{86.6\:\rm N}{5\: \rm kg}=17.3\:\rm m/s^2"